# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

### v1

@@ 3.1.1 Mobility and Transportation Management
Zhou et al. (2021) introduces the Trajectory Generative Adversarial Network (TGAN) to address challenges in identifying and linking human mobility patterns, particularly in location-based social networks (LBSNs). By combining adversarial learning and spatio-temporal data analysis, TGAN significantly enhances trajectory-user linking (TUL) by generating synthetic trajectories that closely mimic real human movement patterns. This approach effectively overcomes issues of data sparsity and imbalanced label distribution in trajectory data, demonstrating superior accuracy in TUL compared to existing methods. The use of GANs for data augmentation in this context marks a substantial advancement in human mobility analysis, offering promising implications for various applications, including urban planning and personalized services. Jilani et al. (2022) presents an innovative approach to traffic congestion analysis using GANs for data augmentation. The research introduces a five-layer convolutional neural network model that leverages GAN-enhanced datasets for improved accuracy in classifying traffic congestion. This method significantly improves identifying congestion patterns, achieving a 98.63% accuracy rate. The study highlights the efficacy of GANs in generating realistic synthetic images, thereby enriching the dataset for more effective traffic management and planning applications. Dabboussi (2023) explores an innovative approach using GANs for data augmentation in traffic management. The study presents a novel GAN-based model that successfully generates synthetic traffic data, addressing the critical issue of data scarcity in Intelligent Transportation Systems (ITS). This method demonstrates a significant improvement in traffic prediction models, enhancing their performance through the addition of realistic, GAN-generated data. The research contributes to developing more intelligent and sustainable transportation systems, underlining the potential of GANs in transforming ITS applications. Islam and Abdel-Aty (2021) presents a novel data augmentation technique using a VAE for improving transportation mode classification with limited training data. The VAE’s decoder is adept at generating synthetic data that closely resembles real smartphone sensor data, effectively enlarging the dataset for machine learning models. This method enhances classification accuracy and addresses the imbalance in datasets by boosting underrepresented classes. Its application in activity recognition using smartphone sensors holds significant potential for vehicle-to-pedestrian (V2P) systems and smart mobility solutions. Islam (2021) presents novel deep learning applications in urban transportation systems. It showcases the use of VAEs and GANs for augmenting data in scenarios with limited datasets. The study highlights the capability of these deep learning models to generate synthetic data, which closely resembles real-world data, thus enhancing model training and predictive performance in transportation applications like crash data analysis and traffic pattern prediction. Islam et al. (2021) introduces a VAE-based data augmentation technique to balance crash and non-crash event datasets in traffic systems. Due to the extreme imbalance in these datasets, traditional machine learning models tend to underperform. The VAE encodes events into a latent space, enabling the generation of synthetic crash data that statistically mirrors real data. This method outperforms traditional oversampling techniques like SMOTE and ADASYN, improving specificity and sensitivity in crash prediction models. It also addresses overfitting issues, showcasing VAE’s capability in generating precise, balanced datasets for effective machine learning applications in traffic safety.
3.1.2 Urban Energy Systems
Data augmentation techniques, aided by generative AI, are increasingly pivotal in building energy modeling and power grid applications. These innovative methods have significantly enhanced the accuracy of predictions and forecasting, thereby streamlining the management of urban energy systems.
Wu et al. (2021) presents a novel approach to forecasting energy consumption using a GAN-enhanced ensemble model. This method combines GANs with ensemble learning, demonstrating significant improvements in forecasting accuracy. The study explores the use of GANs for data augmentation, which effectively enriches the training dataset with high-quality, diverse samples. This approach enhances the model’s robustness and generalization capabilities, showcasing the utility of GANs in improving predictive models in energy management systems. The Boicea et al. (2022) explores using GANs for augmenting data in power load forecasting. It demonstrates GAN’s capability to expand a dataset, enhancing the predictive accuracy of a Convolutional Neural Network model used for load forecasting. This approach addresses the challenge of limited sample size in power load forecasting, showing GANs’ effectiveness in data augmentation and their potential in improving forecasting accuracy in energy management systems. Fan et al. (2022) utilizes deep learning, specifically deep generative modeling, for data augmentation in building energy prediction. This approach significantly improves the accuracy of short-term energy predictions in buildings. The study leverages the capabilities of deep learning to generate synthetic but realistic energy data, addressing the challenge of limited high-quality training data. This innovation demonstrates the potential of deep learning in enhancing data-driven models for smart building operation management, especially in scenarios with data scarcity.
3.1.3 Building and Infrastructure Management
Data augmentation, facilitated by generative AI, significantly enhances the predictive analytics and decision-making processes in building and infrastructure management (Baduge et al., 2022). Its application spans a broad spectrum, from structural detection to maintenance forecasting.
Abdollahi et al. (2020) presents a groundbreaking approach in urban planning by integrating GANs, Genetic Optimization Algorithms (GOAs), and Geographic Information Systems (GIS). This study underscores the capability of deep learning, particularly GANs, in generating complex, high-dimensional urban building images. The innovative use of GANs for data augmentation, combined with the optimization strength of GOAs and the spatial analysis of GIS, demonstrates a novel framework in urban planning, enhancing efficiency, scalability, and decision-making in urban development. Li and Betti (2023) addresses a significant challenge in the Structural Health Monitoring (SHM) of civil infrastructure using machine learning (ML), where a lack of comprehensive datasets for training models under various structural damage conditions is a common issue. To overcome this, the research introduces a novel data augmentation strategy within a vibration-based SHM framework, employing a Conditional Variational Autoencoder (CVAE). The CVAE is used to generate additional data samples of power cepstral coefficients of structural acceleration responses in diverse damage scenarios. This is achieved by training the CVAE to model different statistical distributions of these coefficients, using newly defined conditional variables. Notably, the decoder of the CVAE, once trained, is then used independently for the generation of new samples, enhancing the data augmentation process. Hoeiness et al. (2021) utilizes augmented GANs for predicting wind flow in urban environments. This innovative approach rephrases the problem of 3D flow fields into a 2D image-to-image translation challenge. The study explores the use of GANs, particularly Pix2Pix and CycleGAN, along with U-Net autoencoders, demonstrating GAN’s ability to model complex spatial relationships and generate accurate wind flow predictions. This research represents a significant advancement in urban planning and pedestrian comfort analysis for improving street and infrastructure design.
By generating comprehensive datasets that replicate a vast array of environmental and usage conditions, these AI models empower the development of robust solutions, including innovative designs and optimization strategies. Their applications lead to more resilient and efficient operations and maintenance of buildings and urban infrastructure.
3.2 Data synthesis and Scenario Generation
Synthetic data generation involves artificially creating entirely new data that mimics real-world data and is different from data augmentation, which modifies existing data to create variations for enhancing the size and diversity of data sets for training data models (Figueira and Vaz, 2022). Popular generative AI models like GANs and VAEs are instrumental in synthetic data generation for smart city and urban research, offering diverse, realistic datasets where real data is scarce or sensitive due to privacy issues. In many smart city applications, synthetic data supports scenario generation and analysis, enabling planners to simulate and predict outcomes of future and hypothetical urban developments, traffic patterns, and public services in a controlled and risk-free environment, thereby informing more effective and sustainable urban strategies. This approach leads to a more informed and effective scenario-based planning approach, facilitating the evaluation of potential impacts and outcomes of urban policies or changes before they are implemented in the real world.
3.2.1 Mobility and Transportation Management
@@
In human mobility studies, generative models are widely used to generate synthetic trajectories that can realistically reproduce mobility patterns, as well as to enable predictions of future mobility patterns. Luca et al. (2021) have provided a comprehensive survey on the generative AI applications for creating synthetic trajectories. As examples, Smolyak et al. (2020) innovatively uses GANs for anomaly detection in human mobility data. The authors introduce the Infinite Gaussian Mixture Model (IGMM) coupled with bidirectional GANs (BiGANs), an approach that generates realistic synthetic human mobility data and effectively detects multimodal anomalies. This method marks a significant advancement in anomaly detection, addressing the challenges of multimodal data patterns in human mobility and improving the accuracy of identifying anomalous behaviors or trajectories. (Kosaraju et al., 2019) presents a novel approach to pedestrian trajectory forecasting. It integrates a graph-based GAN with a graph attention network (GAT) and Bicycle-GAN to generate realistic multimodal trajectory predictions as future scenarios. This method effectively models the social interactions of pedestrians and accounts for the multimodal nature of human movement. The research demonstrates significant advancements in trajectory forecasting, contributing to applications like autonomous vehicle navigation and urban planning. (Huang et al., 2019) utilizes a Sequential Variational Autoencoder (SVAE) for reconstructing human mobility trajectories. This approach is novel in its integration of a Variational Autoencoder (VAE) with a sequence-to-sequence model, enabling the efficient capture of salient features in human mobility data. The SVAE model demonstrates the capability of VAEs in approximating complex, high-dimensional distributions like urban mobility patterns. This innovation is crucial for enhancing trajectory reconstruction in urban planning and intelligent transportation systems, addressing challenges of data sparsity and privacy concerns. Bao et al. (2020) introduces an innovative approach using Conditional GANs to estimate human mobility during the COVID-19 pandemic. This research significantly contributes by addressing the challenge of estimating human mobility in rapidly changing social contexts with limited training data. The study employs a domain-constraint correction layer within the GAN’s generator, enhancing the learning process and improving the quality of generated mobility predictions. This method demonstrates the potential of GANs in generating reliable spatiotemporal mobility data under uncertain and dynamically changing conditions, a crucial advancement in predictive modeling during public health crises. Wang et al. (2020) introduces SeqST-GAN, a model utilizing Sequence-to-Sequence (Seq2Seq) Generative Adversarial Networks for predicting urban crowd flow. It uniquely addresses the complex spatial-temporal correlations in urban environments by treating crowd flow data as sequences, akin to frames in a video. The combination of Seq2Seq and GANs enables the model to generate accurate multi-step predictions, overcoming the challenge of blurry predictions often seen in deep learning models. This innovative use of GANs demonstrates their capability in generating sequential, high-dimensional data, marking a significant advancement in urban planning and traffic management.
@@
In many previous efforts that aim to develop intelligent transportation systems for efficient traffic planning and management, generative models are developed to create synthetic data for generating planning and simulation scenarios. As examples, (Wu et al., 2020) explores the use of a Long Short-Term Memory (LSTM) network combined with GAN to generate traffic flow scenarios. This approach leverages the LSTM’s capability in capturing temporal dependencies and GAN’s strength in generating realistic traffic patterns. The innovation lies in accurately modeling and simulating complex traffic dynamics, aiding in traffic management and urban planning. This research highlights the potential of deep learning in enhancing predictive models for urban traffic systems. Demetriou et al. (2023) utilizes GANs to create realistic driving scenarios for autonomous vehicle testing. This research demonstrates the innovative use of GANs in generating complex and diverse driving conditions, aiding in the development and validation of autonomous driving systems. The study highlights GAN’s capability to simulate various road environments and traffic situations, which is crucial for ensuring the safety and reliability of autonomous vehicles in real-world conditions. Mladenov et al. (2022) explores a novel application of GANs in traffic simulation. The study introduces an adversarial variational inference technique for calibrating Origin-Destination (OD) matrices in urban traffic simulators. This method is significant for its ability to mitigate the underdetermination problem in traffic simulation, leveraging GANs to match higher-order moments of field data. This innovative approach enhances the accuracy and robustness of traffic simulations, which is essential for effective urban traffic management and planning.
3.2.2 Urban Analytics
Generative AI models excel in creating realistic, high-dimensional data and data-driven scenarios that can mimic complex urban dynamics and processes based on realistic and historic data. By creating synthetic data and scenarios of various urban growth patterns, demographic distributions, and city layouts, these AI tools provide invaluable insights for urban planning, policy making, and morphology analysis, enabling researchers to explore diverse scenarios and predict future urban transformations effectively (De Sabbata et al., 2023).
In terms of detail applications, Johnsen et al. (2022) explores the use of GANs for synthesizing realistic population datasets. This approach leverages GAN’s capability to generate high-dimensional, diverse data, addressing the challenge of limited real-world demographic data availability. By creating synthetic populations that closely resemble actual demographics, the study showcases GANs’ potential in urban planning and policy making, where accurate population models are crucial. Kempinska and Murcio (2019) explores the use of deep learning, specifically VAEs, in modeling urban street networks. It demonstrates how VAEs can effectively capture key urban network metrics in low-dimensional vectors and generate complex urban forms. This approach, capitalizing on the deep learning capability of generating realistic images, marks significant progress in data-driven urban modeling, leveraging the increasing availability of street network data to enhance the quantitative analysis of urban patterns. Zhang et al. (2022) introduces Metropolitan GAN (MetroGAN), a novel Generative Adversarial Network framework, for simulating urban morphology with geographical knowledge. MetroGAN addresses limitations of previous GAN models such as data sparsity and training instability by incorporating a progressive growing structure and a geographical loss for water area constraints. It significantly outperforms existing urban simulation methods, demonstrating over 20% improvement in all metrics. Interestingly, MetroGAN can generate city shapes using solely physical geography features, showcasing its ability to handle various urban attributes and stabilize GAN-based urban simulation. Allen-Dumas et al. (2022) delves into the application of deep learning, particularly GANs, for analyzing urban morphology. This research highlights the capability of GANs to process and interpret complex urban data, providing insightful analyses of city structures and patterns. The innovative use of GANs in this context offers significant contributions to urban planning and development, enhancing our understanding of urban spaces and their evolution. Sun et al. (2021) explores land-use and land-cover change (LUCC) prediction using deep learning. It introduces a model employing GANs, specifically the pix2pix GAN, combined with an attention mechanism. This approach significantly enhances local spatial detail and diversity in urban planning and LUCC prediction. The model’s unique capability lies in fusing multiscale local spatial information with city planning data to predict future land-use changes, showcasing a novel application of GANs in urban planning and environmental monitoring. Kastner and Dogan (2023) presents an innovative GAN model for predicting urban wind flow. This GAN, designed for image-to-image translation, efficiently maps input urban geometry to predicted wind flow patterns. By utilizing Pix2Pix GAN, integrated with a voxelization process and distance-to-object encoding, the model offers rapid and accurate predictions of urban wind flow. This approach significantly reduces the computational cost and time associated with traditional Computational Fluid Dynamics (CFD) simulations, making it a promising tool for urban design and microclimate analysis.
3.2.3 Urban Energy System

### v2

@@ 3.1.1 Transportation Data Augmentation
Data augmentation refers to techniques used to increase the amount of data by adding slightly modified copies of already existing data or newly created synthetic data from existing data (Lemley et al., 2017). This is particularly common in machine learning and deep learning, where more training data leads to better model performance. Data augmentation is primarily about creating additional data points to improve the robustness and accuracy of models in tasks like image recognition, classification, and predictions (Moreno-Barea et al., 2020). GANs and VAEs are prevalent in smart city applications for data augmentation, given their proficiency in generating high-quality synthetic data. However, generative diffusion models, known for their strengths in denoising and data augmentation, are primarily utilized in medical research rather than urban science. This suggests a potential untapped area for these models in urban applications, where their unique capabilities could offer novel solutions to complex urban challenges. Many previous studies have applied GenAI models to augment transportation data. Zhou et al. (2021) introduces the Trajectory Generative Adversarial Network (TGAN) to address challenges in identifying and linking human mobility patterns, particularly in location-based social networks (LBSNs). By combining adversarial learning and spatio-temporal data analysis, TGAN significantly enhances trajectory-user linking (TUL) by generating synthetic trajectories that closely mimic real human movement patterns. This approach effectively overcomes issues of data sparsity and imbalanced label distribution in trajectory data, demonstrating superior accuracy in TUL compared to existing methods. The use of GANs for data augmentation in this context marks a substantial advancement in human mobility analysis, offering promising implications for various applications, including urban planning and personalized services. Jilani et al. (2022) presents an innovative approach to traffic congestion analysis using GANs for data augmentation. The research introduces a five-layer convolutional neural network model that leverages GAN-enhanced datasets for improved accuracy in classifying traffic congestion. This method significantly improves identifying congestion patterns, achieving a 98.63% accuracy rate. The study highlights the efficacy of GANs in generating realistic synthetic images, thereby enriching the dataset for more effective traffic management and planning applications. Dabboussi (2023) explores an innovative approach using GANs for data augmentation in traffic management. The study presents a novel GAN-based model that successfully generates synthetic traffic data, addressing the critical issue of data scarcity in Intelligent Transportation Systems (ITS). This method demonstrates a significant improvement in traffic prediction models, enhancing their performance through the addition of realistic, GAN-generated data. The research contributes to developing more intelligent and sustainable transportation systems, underlining the potential of GANs in transforming ITS applications. Islam and Abdel-Aty (2021) presents a novel data augmentation technique using a VAE for improving transportation mode classification with limited training data. The VAE’s decoder is adept at generating synthetic data that closely resembles real smartphone sensor data, effectively enlarging the dataset for machine learning models. This method enhances classification accuracy and addresses the imbalance in datasets by boosting underrepresented classes. Its application in activity recognition using smartphone sensors holds significant potential for vehicle-to-pedestrian (V2P) systems and smart mobility solutions. Islam (2021) presents novel deep learning applications in urban transportation systems. It showcases the use of VAEs and GANs for augmenting data in scenarios with limited datasets. The study highlights the capability of these deep learning models to generate synthetic data, which closely resembles real-world data, thus enhancing model training and predictive performance in transportation applications like crash data analysis and traffic pattern prediction. Islam et al. (2021) introduces a VAE-based data augmentation technique to balance crash and non-crash event datasets in traffic systems. Due to the extreme imbalance in these datasets, traditional machine learning models tend to underperform. The VAE encodes events into a latent space, enabling the generation of synthetic crash data that statistically mirrors real data. This method outperforms traditional oversampling techniques like SMOTE and ADASYN, improving specificity and sensitivity in crash prediction models. It also addresses overfitting issues, showcasing VAE’s capability in generating precise, balanced datasets for effective machine learning applications in traffic safety.
3.1.2 Missing Data Imputation
Missing data imputation using GenAI involves utilizing advanced generative models, such as VAEs or GANs, to predict and fill in missing values within a dataset (Shahbazian and Greco, 2023). These models learn the underlying distribution of the complete data and generate plausible values for the missing entries, thereby enhancing the overall quality and integrity of the data (Shammasi, 2024). By capturing complex patterns and relationships in the data, GenAI-based imputation provides more accurate and reliable estimates compared to traditional methods. Among previous studies, GenAI models have been widely used in intelligent transportation systems and smart mobility systems to facilitate the imputation of traffic and mobility data. These models effectively address missing data issues by generating plausible data points, thereby enhancing the accuracy and reliability of data analysis and decision-making processes in urban transportation networks. (Chan et al., 2023). Boquet et al. (2020) introduces a unified model leveraging VAEs to address key challenges in traffic forecasting. The study highlights the use of VAEs for generative modeling, enabling the imputation of missing traffic data, effective dimension reduction, model selection, and anomaly detection. By learning the latent space representation of traffic data, the VAE improves the accuracy and reliability of predictions, compresses data for computational efficiency, and identifies meaningful patterns and anomalies. Extensive experiments on real-world traffic datasets validate the model’s superior performance over traditional methods, demonstrating its potential to enhance intelligent transportation systems. Zhang et al. (2024a) introduces a novel approach utilizing GenAI models to address the challenge of missing traffic data. The study proposes a dynamic multi-level generative adversarial network (MLGAN) model that captures the spatial-temporal dependencies of traffic data through a combination of bidirectional gated recurrent units (BGRU) and graph convolutional networks (GCN). By leveraging these advanced generative models, the MLGAN effectively imputes missing traffic data, even under conditions of high fluctuation and non-uniform distribution of missing values. The research demonstrates that the MLGAN model outperforms traditional imputation methods, providing more accurate and stable imputation results, thereby significantly enhancing the reliability and efficiency of urban traffic management systems. Extensive experiments using real-world data from the PeMS dataset validate the superior performance of the proposed model compared to existing approaches. Many similar GenAI-powered applications used varients of GAN to impute traffic data (Yuan et al., 2022; Kazemi and Meidani, 2021; Yang et al., 2021; Huang et al., 2023a; Chen et al., 2023). A few data have applied GenAI models to imputing missing data in urban mobility systems, such as bike-sharing data (Xiao et al., 2021) and public transit data (Kim et al., 2022).
3.1.3 Generation of Synthetic Mobility Data
In human mobility studies, generative models are widely used to generate synthetic trajectories that can realistically reproduce mobility patterns and enable predictions of future mobility trends. Synthetic data generation involves creating entirely new data points that mimic the statistical properties and structure of the original dataset (Offenhuber, 2024). GenAI models, such as GANs and VAEs, are often used for this purpose. The primary goal is to expand the dataset, providing a larger volume of high-quality data for training machine learning models. This approach is particularly useful when real data is scarce, expensive to collect, or poses privacy concerns. In comparison, the generative task focuses on creating entirely new data points, while data augmentation modifies existing data to produce variations, and data imputation fills in gaps within an existing dataset. Each technique addresses different challenges in data preprocessing and enhancement, contributing to the overall improvement of data quality and utility.
@@ 3.1.3 Generation of Synthetic Mobility Data
As examples, Luca et al. (2021) have provided a comprehensive survey on the GenAI applications for creating synthetic trajectories. As examples, Smolyak et al. (2020) uses GANs for anomaly detection in human mobility data. The authors introduce the Infinite Gaussian Mixture Model (IGMM) coupled with bidirectional GANs (BiGANs), an approach that generates realistic synthetic human mobility data and effectively detects multimodal anomalies. This method marks a significant advancement in anomaly detection, addressing the challenges of multimodal data patterns in human mobility and improving the accuracy of identifying anomalous behaviors or trajectories. (Kosaraju et al., 2019) presents a novel approach to pedestrian trajectory forecasting. It integrates a graph-based GAN with a graph attention network (GAT) and Bicycle-GAN to generate realistic multimodal trajectory predictions as future scenarios. This method effectively models the social interactions of pedestrians and accounts for the multimodal nature of human movement. The research demonstrates significant advancements in trajectory forecasting, contributing to applications like autonomous vehicle navigation and urban planning. (Huang et al., 2019) utilizes a Sequential Variational Autoencoder (SVAE) for reconstructing human mobility trajectories. This approach is novel in its integration of a Variational Autoencoder (VAE) with a sequence-to-sequence model, enabling the efficient capture of salient features in human mobility data. The SVAE model demonstrates the capability of VAEs in approximating complex, high-dimensional distributions like urban mobility patterns. This innovation is crucial for enhancing trajectory reconstruction in urban planning and intelligent transportation systems, addressing challenges of data sparsity and privacy concerns. Bao et al. (2020) introduces an innovative approach using Conditional GANs to estimate human mobility during the COVID-19 pandemic. This research significantly contributes by addressing the challenge of estimating human mobility in rapidly changing social contexts with limited training data. The study employs a domain-constraint correction layer within the GAN’s generator, enhancing the learning process and improving the quality of generated mobility predictions. This method demonstrates the potential of GANs in generating reliable spatiotemporal mobility data under uncertain and dynamically changing conditions, a crucial advancement in predictive modeling during public health crises. Wang et al. (2020a) introduces SeqST-GAN, a model utilizing Sequence-to-Sequence (Seq2Seq) Generative Adversarial Networks for predicting urban crowd flow. It uniquely addresses the complex spatial-temporal correlations in urban environments by treating crowd flow data as sequences, akin to frames in a video. The combination of Seq2Seq and GANs enables the model to generate accurate multi-step predictions, overcoming the challenge of blurry predictions often seen in deep learning models. This innovative use of GANs demonstrates their capability in generating sequential, high-dimensional data, marking a significant advancement in urban planning and traffic management.
3.1.4 Traffic Simulation Scenario Generation
@@
In many previous efforts, generative models have been developed to create synthetic data for planning and simulation scenarios. Numerous studies have specifically applied variants of GANs for these purposes. In terms of simulating traffic flow and dynamics, (Wu et al., 2020) explores a novel approach for generating realistic traffic flow scenarios using GenAI. The study leverages a hybrid model combining Long Short-Term Memory (LSTM) networks with GANs, termed LSTM-GAN. This model captures the temporal dynamics and spatial dependencies of traffic data, enabling the creation of spatiotemporal traffic scenarios for multiple horizons and locations. The GenAI component, specifically the GAN, plays a crucial role in learning from historical traffic data to generate new, realistic traffic flow scenarios. These synthetic scenarios can aid in traffic infrastructure design, road planning, and virtual training environments for intelligent driving systems. The research demonstrates that the LSTM-GAN model outperforms traditional GAN models by better capturing the temporal and spatial correlations inherent in traffic data, thereby providing more accurate and useful synthetic data for intelligent transportation systems. Mladenov et al. (2022) explores a novel application of GANs in traffic simulation. The study introduces an adversarial variational inference technique for calibrating Origin-Destination (OD) matrices in urban traffic simulators. This method is significant for its ability to mitigate the underdetermination problem in traffic simulation, leveraging GANs to match higher-order moments of field data. This innovative approach enhances the accuracy and robustness of traffic simulations, which is essential for effective urban traffic management and planning.
There are also a few similar studies that use GAN-based models to generate traffic simulation scenarios to support autonomous driving research. Demetriou et al. (2023) utilizes GANs to create realistic driving scenarios for autonomous vehicle testing. This research demonstrates the innovative use of GANs in generating complex and diverse driving conditions, aiding in the development and validation of autonomous driving systems. The study highlights GAN’s capability to simulate various road environments and traffic situations, which is crucial for ensuring the safety and reliability of autonomous vehicles in real-world conditions. Li et al. (2021a) introduces a novel approach for generating realistic and diverse traffic scenarios using GenAI. The study presents SceGene, a dynamic traffic scenario generation algorithm inspired by biological intelligence, specifically genetic inheritance and mutation processes. SceGene applies biological processes like crossover, mutation, and natural selection to create new traffic scenarios. This method involves encoding traffic scenario elements into a genotype-like representation, generating new scenarios through genetic operations, and repairing abnormal scenarios using a microscopic driving model. By leveraging these GenAI techniques, SceGene can produce highly realistic and diverse traffic scenarios in an interpretable and controllable way, significantly improving the efficiency and coverage of simulation-based autonomous driving tests. The generated scenarios enhance the robustness and safety of autonomous driving algorithms by providing a comprehensive range of testing conditions. Apart from GAN-based applications, Kempinska and Murcio (2019) explores the use of deep learning, specifically VAEs, in modeling urban street networks. It demonstrates how VAEs can effectively capture key urban network metrics in low-dimensional vectors and generate complex urban forms. This approach, capitalizing on the deep learning capability of generating realistic images, marks significant progress in data-driven urban mobility modeling, leveraging the increasing availability of street network data to enhance the quantitative analysis of urban patterns.
With the recent advancement in LLMs, a few emerging studies explores the use of GPT to generate scenarios. As a exmaple, Güzay et al. (2023) explores the innovative use of Large Language Models (LLMs) like GPT-4 to generate traffic scenarios for simulation purposes. The study highlights how LLMs, typically used in natural language processing, can be adapted to create XML-based simulation files for tools like SUMO (Simulation of Urban Mobility). By leveraging the generative capabilities of LLMs, the researchers aim to simplify the scenario creation process, making it more fluent and flexible compared to traditional methods. The approach allows users to describe traffic scenarios linguistically, which are then processed by the model to produce accurate simulation files. This application of GenAI not only streamlines the development of traffic scenarios for safety testing, autonomous vehicle training, and regulatory compliance but also addresses challenges related to the complexity and rigidity of existing simulation setup methods.
Apart from the four major types of GenAI models, our literature search has also identified the application of other types of GenAI models. Tan et al. (2021) presents a novel approach to generating realistic traffic scenarios using GenAI. The study introduces SceneGen, a neural autoregressive model that generates traffic scenes by sequentially inserting actors such as vehicles, pedestrians, and bicyclists into a given environment based on the ego-vehicle’s state and a high-definition map. This method avoids the need for hand-crafted rules and heuristics, allowing for more complex and diverse traffic scenarios that better mimic real-world conditions. The GenAI model used in SceneGen captures the spatial and temporal dependencies of traffic scenes, significantly improving the fidelity of the simulations. The results demonstrate that SceneGen can generate realistic traffic scenarios that enhance the training and testing of self-driving vehicles, providing more accurate and scalable solutions compared to traditional methods.
3.2 Urban Energy Systems
Among the 479 selected articles refined for their relevance to urban energy systems, we conducted a further screening based on their titles and abstracts to summarize the major GenAI use cases that can be incorporated into urban digital twins in the context of urban energy management and grid operation. These use cases are discussed in the following subsections.
3.2.1 Energy Data Enhancement
Data enhancement encompasses various techniques aimed at improving the quality, quantity, and diversity of data. Among the selected studies, two frequently used GenAI-powered techniques are data augmentation and missing data imputation. While some autoencoder-based GenAI models, such as VAEs, are capable of performing other data enhancement tasks, such as dimensionality reduction and anomaly detection, these features do not leverage the generative capabilities of the AI. Consequently, studies related to these features are not discussed in this section.
Data augmentation techniques, powered by GenAI, are becoming increasingly pivotal in building energy modeling and power grid applications. These innovative methods have substantially improved the accuracy of predictions and forecasts, thereby optimizing the management of urban energy systems. Wu et al. (2021) presents a novel approach to forecasting energy consumption using a GAN-enhanced ensemble model. This method combines GANs with ensemble learning, demonstrating significant improvements in forecasting accuracy. The study explores the use of GANs for data augmentation, which effectively enriches the training dataset with high-quality, diverse samples. This approach enhances the model’s robustness and generalization capabilities, showcasing the utility of GANs in improving predictive models in energy management systems. The Boicea et al. (2022) explores using GANs for augmenting data in power load forecasting. It demonstrates GAN’s capability to expand a dataset, enhancing the predictive accuracy of a Convolutional Neural Network model used for load forecasting. This approach addresses the challenge of limited sample size in power load forecasting, showing GANs’ effectiveness in data augmentation and their potential in improving forecasting accuracy in energy management systems. Fan et al. (2022b) utilizes deep learning, specifically deep generative modeling, for data augmentation in building energy prediction. This approach significantly improves the accuracy of short-term energy predictions in buildings. The study leverages the capabilities of deep learning to generate synthetic but realistic energy data, addressing the challenge of limited high-quality training data. This innovation demonstrates the potential of deep learning in enhancing data-driven models for smart building operation management, especially in scenarios with data scarcity.
A few studies have utilize GenAI models to facilitate missing data imputation. Hwang and Suh (2024) proposes a novel method for addressing missing data in electricity consumption datasets, crucial for effective energy management in smart grid environments. The study introduces the CC-GAIN model, which integrates clustering, classification, and GANs to impute missing data. This model leverages time-series and pattern features of electricity consumption data to enhance imputation accuracy. The CC-GAIN model significantly outperforms existing imputation techniques by effectively handling diverse types and rates of missing data, making it a robust solution for maintaining high-quality data essential for smart grid operations. The research highlights the model’s superior performance through extensive evaluations and ablation studies, demonstrating its potential to improve data reliability and support sustainable energy management. Ryu et al. (2020) explores a deep learning approach for addressing missing data in smart meter datasets. The study introduces a framework using a Denoising Autoencoder (DAE) to impute missing electricity consumption values. This GenAI technique is compared with traditional imputation methods like linear interpolation and historical averaging, as well as other generative models such as the VAE and Wasserstein Autoencoder (WAE). The DAE-based approach significantly outperforms these methods, showing up to 28.9% lower point-wise errors and up to 56% lower daily-accumulated errors. The paper highlights the robustness and accuracy of the DAE model in reconstructing incomplete daily load profiles, thus enhancing the reliability of smart grid data analytics.
3.2.2 Data Synthesis and Scenario Generation